10 kg of water is placed in an enclosed volume of 3m³.Heat is added until the temperature is 250°C.For the final state; A) the pressure is? B) the mass of vapour is? C) the enthalpy of the vapour is?
1
Expert's answer
2020-09-22T16:11:30-0400

a) We use the table of saturated water - temperature . In the quality region the pressure as "p = 3976.2 kPa"

b) To find the mass of the vapor we must determine the quality. We use the next equation-

"U = Uf + x(Ug - Uf)"

Then,

"0.3 m3\/kg = 0.00125 m3\/kg + x( 0.05008 - 0.00125 ) m3\/kg"

"0.3 = 0.00125 + 0.04883x"

"0.04883x = 0.3 - 0.00125"

"0.04883x = 0.29875"

"x = 0.29875 (m3\/kg) \/ 0.04883 (m3\/kg)"

"x = 6.11816"

Using the relationship of-

"X = mg \/ m ,"

we find the vapor mass

"mg = xm = 6.11816 \times10 kg = 61.1816 kg"

To find the enthalpy of the vapor-

Using the relationships

"dinP = -\frac{d(AHvap)}{R}\frac{1}{T}"

"\Delta Нvap = (\frac{1}{3976.2}) \times (523?) \times8.314"

"\Delta Нvap = 571.93Ј"